### v1

# AI Coders Are Among Us: Rethinking Programming Language Grammar Towards Efficient Code Generation

Abstract.
Besides humans and machines, Artificial Intelligence (AI) models have emerged to be another important audience of programming languages, as we come to the era of large language models (LLMs). LLMs can now excel at coding competitions and even program like developers to address various tasks, such as math calculation. Yet, the grammar and layout of existing programs are designed for humans. Particularly, abundant grammar tokens and formatting tokens are included to make the code more readable to humans. While beneficial, such a human-centric design imposes an unnecessary computational burden on LLMs where each token, either consumed or generated, consumes computational resources.
To improve inference efficiency and reduce computational costs, we propose the concept of AI-oriented grammar, which aims to represent the code in a way that better suits the working mechanism of AI models. Code written with AI-oriented grammar discards formats and uses a minimum number of tokens to convey code semantics effectively. To demonstrate the feasibility of this concept, we explore and implement the first AI-oriented grammar for Python, named Simple Python (SimPy). SimPy is crafted by revising the original Python grammar through a series of heuristic rules. Programs written in SimPy maintain identical Abstract Syntax Tree (AST) structures to those in standard Python, allowing execution via a modified AST parser. In addition, we explore methods to enable existing LLMs to proficiently understand and use SimPy, and ensure the changes remain imperceptible for human developers. Compared with the original Python, SimPy not only reduces token usage by 13.5% and 10.4% for CodeLlama and GPT-4, but can also achieve equivalent, even improved, performance over the models trained on Python code. With these promising results, we encourage further contributions to the development of AI-oriented grammar within our community.
conference: ; ;
1. Introduction
High-level programming languages, such as the widely adopted Python, are designed with two types of audiences in mind (Casalnuovo et al., 2020): machines that compile and execute programs and humans who read, comprehend, and write programs. Machines focus on the operational semantics of programs, while humans additionally emphasize programs’ readability, a key factor in understanding source code. For example, one of the guiding design principles for Python (Peters, 2023) is that “readability counts.” Consequently, these languages incorporate numerous human-centric design elements within their grammar. For example, programming languages utilize explicit delimiters to separate code structures. These elements, while enhancing human readability, may not be necessary to convey the program’s operational semantics.
Recently, the audiences for programming languages have expanded to include AI models, particularly Large Language Models (LLMs) that can analyze, generate, and execute code. This is evident by the impressive performance that LLMs achieved in code generation (Hou et al., 2023). For example, a recently released LLM, AlphaCode2 (AlphaCode Team, 2023), is reported to perform better than 85% of human participants in a programming competition. Moreover, many LLM-powered assistants, such as ChatGPT (OpenAI, 2023a) and Bard (Google, 2023), are now equipped with code execution environments, which enable them to execute generated code and provide responses based on the results. Thus, the role of LLMs has evolved from mere code generators to active “developers” that utilize programming to accomplish various tasks, including mathematical computations and file processing. This paradigm shift indicates a new era in which AI models emerge as an important group of users of programming languages.
Figure 1. An illustration of how LLMs and human programmers perceive the source code.
While AI models have taken on the new role, the design of code grammar is yet to evolve to accommodate their unique needs. The readability of the source code could impose an additional computational burden on LLMs to read and generate programs, as extra tokens are particularly placed to make the code layout more friendly for reading. However, readability may not be essential for LLMs to perform coding tasks. Studies have revealed that code models do not capture much information relevant to readability (Troshin and Chirkova, 2022), and readability-enhancing symbols like “:” received significantly lower attention compared to other elements such as variable names(Zeng et al., 2022). We illustrate how a program is perceived by humans and AI models in Figure 1. When certain elements that enhance readability are omitted from the code while preserving its underlying semantics, it becomes more challenging for humans to interpret, yet AI models can process the code more efficiently. The observation makes us wonder: What is a suitable grammar for AI models? Exploring this question is vital for optimizing the efficiency of LLMs and reducing energy waste in dealing with unnecessary tokens, especially given that the high operational cost of LLMs sets a big challenge for providers to generate profit (https://www.wsj.com/tech/ai/ais-costly-buildup-could-make-early-products-a-hard-sell-bdd29b9f, [n. d.]) from them. As AI models consume and generate source code in a token-by-token manner, with one feed-forward process for each token, reducing the tokens in code representation holds the potential to proportionally reduce the time and energy cost.
This motivates us to propose the concept of AI-Oriented Grammar, a grammar specifically designed for AI models instead of humans. The core idea is to derive grammar rules that keep the code representations concise (with a minimal number of tokens to AI models). Notably, the code crafted in this grammar can be parsed with its adapted parser and then executed to obtain the same result as the original grammar. A few challenges are in the way of designing such a new grammar and melting it into AI models. The AI models are expected to not only comprehend code written in this grammar but also generate code following its rules to better serve the goal of efficiency. At the same time, human developers, who are directing the development, expect to work with grammar that they find friendly and are familiar with.
Given these challenges, the realization of this concept remains uncertain. To assess the feasibility of AI-oriented grammar, we embarked on an exploratory study. The study seeks to consider the implications and limitations of integrating AI-oriented grammar into the existing code generation workflow. It is guided by three research questions, each addressing a key challenge.
RQ1. What is the token reduction capacity of AI-oriented grammar in source code?
Whether and to what extent an AI-oriented grammar can reduce the tokens remains an open question. We fill this gap by implementing a proof-of-concept AI-oriented grammar and assessing its performance. Specifically, we explore a new grammar for Python, named SimPy, by heuristically modifying the standard grammar. Compared to the standard Python grammar, we prohibit using tokens popularly hired to style the code appearance, e.g., whitespace and newline, and simplify keywords, operators, and delimiters to a more compact form. The modifications are designed to be simple, as this is the first attempt to explore such AI-oriented grammar. We also developed an AST parser for SimPy that can parse its code into the same AST as standard Python code, as well as a converter for seamless code transitions between SimPy and Python code. A comparative analysis of SimPy and Python’s original grammar was conducted using tens of tokenizers employed by existing LLMs. The findings indicate a notable reduction in token usage when employing SimPy, with decreases ranging between 8.6% and 34.7%, thus reducing the time and computational cost during inference by a similar level (Kaplan et al., 2020). For example, the tokenizer of GPT-4 demonstrates a significantly enhanced efficiency with SimPy, achieving a 10.4% reduction in token size.
RQ2. How can AI models understand AI-oriented grammar?
Prior research demonstrates that AI models are capable of comprehending human-centric grammars of existing programming languages (Hou et al., 2023). However, how these models can learn AI-oriented grammar remains unexplored. We thus further experiment with SimPy to find an effective way. We explored two different training strategies: directly training a model on a SimPy-based code dataset (converted seamlessly from a Python dataset) and fine-tuning a model, originally trained with a Python dataset, on the SimPy dataset. A control group, where a model is directly trained on the Python code dataset, is also included for comparison. The models trained with either strategy should achieve at least equivalent accuracy compared with the control group. Otherwise, it would be impractical to adopt AI-oriented grammar. For each training strategy, we experiment with three models, CodeGen-NL, TinyLlama and Pythia. The experimental results reveal that models initially trained with Python can adapt effectively to SimPy. For instance, our CodeGen model, initially trained on Python, attains a 7.32% Pass@10 on HumanEval, while it even reaches a 9.15% Pass@10 after the fine-tuning on SimPy.
RQ3. How can AI-oriented grammar support real-world scenarios?
Given that AI-oriented grammar may compromise human readability, its application is somewhat restricted. Thus, a remaining challenge for AI-oriented grammar is: how could an AI-oriented grammar be used in real-world scenarios, particularly when human-readable source code is necessary? To address this challenge, we first discuss the basic usage scenario of AI-oriented grammar, i.e., the scenarios where the code generated by the AI models is not intended to be displayed to human users. In such scenarios, such as AI agents (Significant Gravitas, [n. d.]) for ordinary users who are unfamiliar with programming, the model just generates code in AI-oriented grammar, executes it, and provides outputs to the user based on execution results. However, there are still many scenarios where human developers need to review the code, such as coding assistants. We thus propose an inference framework for code generation named DualCode. DualCode utilizes a rule-based converter to convert code between these grammars, ensuring that users interact with human-readable code as usual, while the model benefits from the efficiency of AI-oriented grammar. Our tests confirm that DualCode introduces negligible latency, with the converter of SimPy processing code under 500 tokens in less than 1.0 ms.
We will release the source code once the paper is accepted. The contributions of this paper are summarized as follows:
  • We propose the concept of AI-oriented grammar and empirically explore its feasibility and potential, paving the way for future improvements in programming language design that prioritize AI efficiency.
  • We implement the first AI-oriented grammar for Python, named SimPy, which can reduce at least 8.3% tokens in Python source code.
  • We propose a novel code generation framework, DualCode, expanding the applicability of AI-oriented grammar beyond AI-only scenarios with negligible additional latency.
2. Motivation
In this section, we critically examine the human-centric elements in the grammar of current programming languages and propose the concept of AI-oriented grammar. Following this, we introduce the dataset an empirical study driven by three research questions to explore the feasibility of this concept.
2.1. Human-centric Grammar Design
Figure 2. Illustrations of human-centric design elements in Python.
As discussed in Section 1, modern programming languages are predominantly designed with human-centric grammar. This design philosophy originates from the longstanding reality that humans were the only developers for decades. In the current era of LLMs, this human-centric design philosophy has not been significantly challenged. To better understand this orientation, we critically examine the grammar of widely used programming languages, focusing on lexical and syntactical elements that enhance human readability. Below, we summarize the identified patterns and provide examples in Figure 2:
Visual Coding Style The programming language grammar is deliberately crafted to accommodate diverse coding styles. Although not mandatory, styles like those recommended in the Python PEP8 guide (van Rossum et al., 2023) rely on grammatical support. For example, the coding style requires the programs to be written in multiple lines instead of a single extremely long line, easing human code review on screens. This necessitates several lexical elements: line breaks to separate lines, indents to visualize code blocks, and line continuation characters for splitting long lines. Figure 2 demonstrates these aspects, with line breaks and indents highlighted in purple. Similarly, the coding style suggests surrounding each binary operator with a single white space on either side. Therefore, lexical grammar must accommodate such stylistic elements, even if they may not contribute to the core semantics in parsing.
Intuitive Notations The human-centric syntax of programming languages is designed to be intuitively understandable to humans. Common operators like “+” for addition and “=” for assignment are chosen for their familiarity, and derivations like the augmented assignment operator “+=” maintain this intuitive connection. Although potentially more concise symbols could replace these (e.g., using a brand-new symbol “\$” for “+=”), they are still deliberately designed to maintain human readability. Similarly, for structural clarity, programming languages often employ explicit delimiters, such as symbols or keywords, to define code structures despite these delimiters not being essential for parsing. For instance, Python’s compound statements, such as the if statement and for statement, use a colon to demarcate the header from the body. While a parser might deduce these components from line breaks alone, the colon acts as a visual aid, as illustrated in Figure 2 where colons are highlighted in red. This emphasis on intuitive notation and explicit delimiters, although not essential for parsing, significantly aids human comprehension.
2.2. AI-Oriented Grammar
Grammar is a rule set that defines how the source code should describe the programming language’s semantics in aspects of lexis and syntax, using notations such as symbols and keywords. The primary function of the notations in the grammar is two-fold: to define a program’s structure for machine execution and to enhance visual comprehension for human readability. Given that AI models do not require assistance in visual comprehension, the focus of AI-oriented grammar is solely on structural definition. We thus consider a notation unnecessary for AI models if it does not contribute to accurate parsing by the parser. AI-oriented grammar is designed with indispensable notations.
In the design process of a programming language, semantics are defined first, followed by the development of a grammar to represent them. Therefore, employing AI-oriented grammar does not alter the fundamental semantics of the programming language. Codes written in either AI-oriented or human-centric grammar of a programming language can still be parsed into the same AST and then executed.
2.3. Python Code Dataset for Our Study
As a newly proposed concept, we are still unclear whether AI-oriented grammar can be realized and what scenarios it can be applied to. To address these uncertainties and explore the potential of AI-oriented grammar, we conduct an empirical study guided by three critical research questions, respectively introduced in Section 3Section 4Section 5. Our study is centered around Python, the main programming language of the execution environment for LLMs like GPT-4 and Bard to address programming-required tasks. We utilize the Python subset of starcoderdata (Li et al., 2023), a filtered variant of The Stack dataset (Kocetkov et al., 2022), a comprehensive collection of over 20 million code files sourced from open-source GitHub repositories. We keep the code files from the repositories with over 100 stars, resulting in 623,887 code files. The dataset is partitioned into training and validation sets in a 95:5 ratio. We do not create a separate testing set, as we plan to evaluate the model’s performance using other established evaluation datasets. The code snippets in the evaluation datasets are excluded from the training dataset.
3. Token Reduction with AI-oriented Grammar (RQ1)
In this section, we present an instance of AI-oriented grammar to answer RQ1: What is the token reduction capacity of AI-oriented grammar in source code? We propose an AI-oriented grammar for Python as a proof-of-concept (Section 3.1) and then proceed to evaluate the extent of token reduction achievable with this grammar (Section 3.2).
Figure 3. A comparison between Python and SimPy source code, tokenized by GPT-4’s tokenizer. Continuous characters with the same background color represent the same token. Notably, there are no line breaks in SimPy and we add these line breaks in the figure for our human readers.
Table 1. Comparison of grammar specifications for Python and SimPy, using the official Python grammar notation ((Python, 2023)). The grammar tokens are in blue. The table also includes the count of grammar tokens for each production: “N” represents the number of lines, “n” signifies the count of repetitive elements, and “?” indicates that the number of tokens is conditional.
{tblr}
row1 = c, colsep = 4pt, cell21 = r=2c, cell22 = c, cell23 = fg=blue, cell24 = c, cell32 = c, cell33 = fg=blue, cell34 = c, cell41 = r=2c, cell42 = c, cell44 = c, cell52 = c, cell54 = c, cell61 = r=2c, cell62 = c, cell64 = c, cell72 = c, cell74 = c, cell81 = r=2c, cell82 = c, cell84 = c, cell92 = c, cell94 = c, cell101 = r=2c, cell102 = c, cell104 = c, cell112 = c, cell114 = c, cell121 = r=2c, cell122 = c, cell124 = c, cell132 = c, cell134 = c, cell141 = r=2c, cell142 = c, cell144 = c, cell152 = c, cell154 = c, cell161 = r=2c, cell162 = c, cell164 = c, cell172 = c, cell174 = c, cell181 = r=2c, cell182 = c, cell184 = c, cell192 = c, cell194 = c, cell201 = r=2c, cell202 = c, cell204 = c, cell212 = c, cell214 = c, vlines, hline1-2,4,6,8,10,12,14,16,18,20,22 = -, Name & Grammar Specification #Token
block Python NEWLINE INDENT statements DEDENT  N+1
SimPy ‘¡block_start¿’ statements ‘¡block_end¿’ 2
function_def Python ‘def’ NAME [type_params] ‘(’ [params] ‘)’ [‘-¿’ expression ] ‘:’ [func_type_comment] block 4+1?
SimPy ‘¡def_stmt¿’ NAME [type_params] [params] [‘¡arrow¿’ expression ] [func_type_comment] block 1+1?
class_def Python  ‘class’ NAME [‘(’ [arguments] ‘)’ ] ‘:’ block  2+2?
SimPy  ‘¡class_stmt¿’ NAME [‘(’ [arguments] ‘)’ ]  block  1+2?
if_stmt Python  ‘if’ named_expression ‘:’ block elif_stmt  2
SimPy  ‘¡if_stmt¿’ named_expression block elif_stmt  1
for_stmt Python  ‘for’ star_targets ‘in’ ~ star_expressions ‘:’ [TYPE_COMMENT] block [else_block]  3
SimPy  ‘¡for_stmt¿’ star_targets ~ star_expressions [TYPE_COMMENT] block [else_block]  1
with_stmt Python  ‘with’ ‘,’.with_item+ ‘:’ [TYPE_COMMENT] block  2+n
SimPy  ‘¡with_stmt¿’ ‘ ’.with_item+ [TYPE_COMMENT] block  1
try_stmt Python  ‘try’ ‘:’ block except_block+ [else_block] [finally_block]  2
SimPy  ‘¡try_stmt¿’ block except_block+ [else_block] [finally_block]  1
while_stmt Python  ‘while’ named_expression ‘:’ block [else_block]  2
SimPy  ‘¡while_stmt¿’ named_expression block [else_block]  1
import_from Python  ‘from’ (‘.’ | ‘…’)* dotted_name ‘import’ import_from_targets 2+n?
SimPy  ‘¡from_import_stmt¿’ (‘.’ | ‘…’)* dotted_name import_from_targets 1+n?
simple_stmts Python  ‘;’.simple_stmt+ [‘;’] NEWLINE n+1+1?
SimPy  [‘¡line_sep¿’].simple_stmt+ [‘¡line_sep¿’] n?+1?
3.1. An AI-oriented grammar for Python
To demonstrate the potential benefits of AI-oriented grammar, we propose and implement the first AI-oriented grammar for Python, named Simple Python (SimPy), as a proof-of-concept. SimPy is derived from the original grammar of Python, where we simplify the grammar specification according to the philosophy of AI-oriented grammar. Alongside SimPy, we develop a toolkit including a parser to interpret SimPy source code into Python’s AST, and a converter for seamless code translation between SimPy and Python.
3.1.1. Design
The design of SimPy is guided by the philosophy of AI-oriented grammar, resulting in a grammar that prioritizes AI processing efficiency over human readability. When being tokenized, the code of SimPy is considerably more compact than Python’s, facilitating faster model processing. A comparative illustration of SimPy and Python code, showcasing the same AST but with fewer tokens in SimPy (measured by the tokenizer of GPT-4), is presented in Figure 3. Next, we introduce what is changed to craft SimPy.
The modifications to production rules are implemented for the grammar specification of Python 3.12 (the latest version). The modifications for SimPy are driven by our understanding of AI models and the semantics of Python, which is not guaranteed to be optimal but is sufficient to serve as a proof-of-concept demonstration for AI-oriented grammar. Limited by the space, we will only introduce the major modifications here, with the complete grammar specification available in our artifact. Below, we highlight major categories of changes implemented:
Replace notations with tokens. In SimPy, we replace keywords and most symbols (e.g., “in”, “¿”, “¿=”) with distinct token placeholders (e.g., “¡in¿”, “¡g¿”, “¡ge¿”). These placeholders require the corporation from the tokenizer side and should be recognized as individual tokens or as part of a token. This approach results in the creation of 78 new tokens. Some single-character symbols, like “.”, “:”, “=”, are retained as they are unambiguous and represent atomic units already optimized for tokenization. This replacement goes beyond a simple tokenization. Firstly, it eliminates the need for surrounding whitespace, as seen in the conversion from Python’s “a in b” to SimPy’s “a¡in¿b”. Secondly, it differentiates the same notation used in different contexts, such as “¡if_stmt¿” and “¡if¿”, to prevent potential conflicts during the creation of SimPy. Finally, it consolidates multi-character symbols in Python, which may be recognized as multiple tokens, into single tokens in SimPy, ensuring the integrity of their meaning.
[…]
Simplify grammar tokens. For every grammar token in every production, we review whether it can be removed, merged with others, or replaced with white spaces. Using white space aligns with the nature of current tokenizers, which usually treat white spaces as part of a token to reduce the token number. In Table 1, we compare the grammar specifications of key productions before and after such simplification and calculate the resulting token reductions. For example, the grammar tokens in the “function_def” statement are reduced from 4 of 1, with an additional token being optional. Our goal is to minimize grammar token usage in every production. However, certain design decisions in one production might introduce parsing ambiguities when considering the grammar globally. For example, in Python, strings can be concatenated without separators, such as “‘hello’ ‘world’ ”. It causes a conflict when we no longer use commas to separate the elements in a list, such as “[‘1’ ‘2’ ‘3’]”, causing ambiguity about whether “‘1’ ‘2’ ‘3”’ is a single concatenated string or three separate strings. To address such ambiguities, strategic trade-offs are necessary. For example, we add an additional grammar token, “¡concat¿”, between the juxtaposed strings so that we can eliminate the commas in a list. Though it leads to an increase in the token number for the juxtaposed strings, it is still beneficial for the overall efficiency of SimPy.
3.1.2. Unambiguity of SimPy
To determine whether a grammar has ambiguity is theoretically undecidable (Ginsburg and Ullian, 1966). In practice, parser generator tools are commonly hired to check for ambiguities in grammar, including those of popular programming languages (Grune and Jacobs, 2007). A parser generator can find a wide range of ambiguities in the grammar, such as conflicts that arise when the parser has two possible actions at one step. Practically, this is almost the best way to check the ambiguity of SimPy. We have successfully generated parsers for SimPy using the GLR (generalized left-to-right rightmost derivation parser) parsing algorithm (Lang, 1974) form tree-sitter (tree sitter, 2023), where no ambiguity is detected.
Next, we provide an analytical discussion about why our transformations are unlikely to introduce ambiguity to the grammar. First of all, the transformations are only made to terminal notations, which act as keywords or delimiters. Changes made to keywords are guaranteed to represent its unique semantics, while changes made to delimiters should not affect the recognition of the construct, and its precedent and subsequent constructs.
Case I: New unique notations are added or introduced as replacements.
Importantly, different notations are not replaced with the same new notations. To this end, the new notations do not interfere with production rules for which the transformation is not applicable. Given that they are semantically equivalent notations as the original one, the parsing of the affected production rules remains the same. For example, replacing the ‘NEWLINE INDENT’ in the production rule of block (see Table 1) with ‘¡blcok_start¿’ conveys the same semantics that a block is about to start.
Case II: Existing notations are removed.
Arbitrary removal notations may introduce ambiguity to the grammar. We carefully design a few heuristics when removing notations such that they are unlikely to cause problems.
  • Remove notations that hold redundant semantics as their adjacent notations. For example, ‘:’ in many statements indicates the end of the previous construct and the start of a new construct, e.g., in ‘if’ named_expression ‘:’ block elif_stmt. However, the block construct initiates with its own starting symbol, making the construct itself distinguishable from any previous construct. Hence, removing ‘:’ is safe for this case.
  • Remove delimiters used to scope a construct when the scope of its precedent and subsequent constructs are clear. For example, the ‘(’ and ‘)’ for parameters are actually unnecessary in function_def_raw := ‘def’ NAME [type_params] ‘(’ [params] ‘)’ [‘-¿’ expression ] ‘:’ [func_type_comment] block. NAME is an atomic token, thus will not interfere the beginning of parameters when type_params is absent. type_params are surrounded by ‘[’ and ‘]’, making their presence not an issue for recognizing params. Hence, ‘(’ can be safely removed. Now, looking at the subsequent constructs, [‘-¿’ expression ], ‘:’, [func_type_comment], or block possesses a unique indicator of their beginning. Hence, ‘)’ can be safely removed as well. Another example is the ‘import’ keyword in import_from := ‘from’ (‘.’ — ‘…’)* dotted_name ‘import’ import_from_targets. Since dotted_name is a must and contains no white spaces, hence the white space between dotted_name and import_from_targets can perfectly separate these two constructs. Removing ‘import’ is also fine.
3.1.3. Semantic equivalence between SimPy and Python
SimPy is designed as a simplified grammar of Python, which means a program written in Python can be equivalently and deterministically transformed to its counterpart in SimPy , and vice versa. In other words, Python and SimPy are semantically equivalent. We prove this statement in Theorem 2.
Formally, we define a grammar $G$ and a grammar $G^{\prime}$. $G^{\prime}$ is obtained via a transformation $T$ to the production rules in $G$. Given a production rule, $T$ is restricted to adding, replacing, or removing a terminal notation or a sequence of terminal notations. The transformation between Python and SimPy is an instance complying with this restriction. For example, $T$(block := NEWLINE INDENT statements DEDENT ) = block:= ‘¡block_start¿’ statements ‘¡block_end¿’. The transformations can be easily extended to parse trees, where $T$ transforms the (production rule) nodes of a tree. Given a parse tree $t$, we define its corresponding program as $p=\tau(t)$. We first prove that the parse tree of a program written in $G$ is equivalent to the parse tree of its counterpart in $G^{\prime}$ modulo the transformation $T$ in Lemma 1. In Theorem 2, we further prove that these program pairs share the same ASTs, and naturally the same semantics.
Lemma 0.
Given any program $p$, the parse tree $t$ generated by grammar $G$ is equivalent to the parse tree $t^{\prime}$ generated from $p^{\prime}=\tau(T(t))$ by grammar $G^{\prime}$ modulo the transformation $T$, if no ambiguity exists in $G^{\prime}$. That is $t^{\prime}=T(t)$.
Proof.
We give the proof by structural induction on $p$.
Base case: $p$ is an atomic program construct. This case is trivial as we don’t have any transformation for the atomic program. Naturally, the parse trees are identical.
Inductive case: We present the proof based on different compound language constructs.
Take the block construct of SimPy for an example. Assuming $p$ = NEWLINE INDENT statements DEDENT, the parse tree for $p$ is $t$=block (NEWLINE, INDENT, statements($t_{0}$), DEDENT), where $t_{0}$ is the parse tree for the stmt construct.
By definition, $p^{\prime}$ = $T(p)$ = ‘¡block_start¿’ $T$(statements) ‘¡block_end¿’.
Obviously, $p^{\prime}$ still matches the production rule of block in $G^{\prime}$. Given no ambiguity in $G^{\prime}$, its parse tree is deterministic. Thus, we have $t^{\prime}$ = block(‘¡block_start¿’, statements($t_{0}^{\prime}$), ‘¡block_end¿’), where the parse tree for the statements construct is $t_{0}^{\prime}$.
By the induction hypothesis, $T(t_{0})=t_{0}^{\prime}$. Applying $T$ to $t$, we have $T(t)=t^{\prime}$.
Similarly, we can prove other constructs.
Hence, the lemma is proved.
Theorem 2.
Python and SimPy are semantically equivalent.
Proof.
Two programs are semantically equivalent if they share the same AST. In the next, we prove that for any program $p$ in Python, it shares the same AST as its counterpart $p^{\prime}$ in SimPy , and vice versa.
We assume the transformation from Python to SimPy is $T$. Assuming the parse tree for $p$ is $t$ and the parse tree for $p^{\prime}$ is $t^{\prime}$, with Lemma 1, we know that $t^{\prime}=T(t).$ ASTs are obtained from parse trees by abstracting away some concrete syntax details but retaining the essential semantic structures. Since $T$ is a non-semantic transformation, $t$ and $t^{\prime}$ differ only at the non-semantic nodes, such that the parse trees are processed to the same AST. Hence, we can prove that for any program in Python, its counterpart in SimPy is semantically equivalent to it. Similarly, we can prove that for any program in SimPy , its counterpart in Python is semantically equivalent to it as well. Thus, the theorem is proved.
3.1.4. Implementation
Based on the grammar specifications of SimPy, we develop a toolkit for it, including an AST parser for SimPy code and a converter for seamless translation between SimPy and Python source codes. The parser is built upon tree-sitter (tree sitter, 2023), a popular parser generator tool. We first describe the grammar specification of SimPy in the configuration file of the tree-sitter and then generate the parser. With the help of the GLR algorithm from the tree-sitter, we ensure SimPy resolves all the conflicts and no ambiguity exists. The generated parser can parse the SimPy source code into the AST of Python. Based on this parser, we further implement a converter, where specific conversion rules are established for each node of the AST. From a pragmatic point of view, we test our implemented toolkits by conducting round-trip transformations, where Python source code is first converted into SimPy code and subsequently retranslated back to Python. Our first tests on the Python dataset revealed that, ignoring all whitespace, the textual content of the code remains unchanged after the transformation. In addition, we assess its soundness through execution results. We perform the round-trip transformation to the ground-truth code snippets of HumanEval and run the test cases on both the transformed and the original code. The execution results of all the transformed code and the original code are exactly the same, which also indicates the soundness of our implementation.
3.2. Experiments of RQ1
In this section, we detail the tokenizers employed in our experiments and describe the experimental methodology and its corresponding results.
3.2.1. Tokenizers
Our experiments encompass a broad spectrum of tokenizers from various LLMs. The main difference between them is the training corpus, leading to different token vocabularies.
GPT-2 (Radford et al., 2019), Codex (Chen et al., 2021), GPT-3.5 (OpenAI, 2023b), GPT-4 (OpenAI, 2023c): These tokenizers, released by OpenAI, are trained on a mixed corpus, including both natural language and programming language, with GPT-4 being the latest version offering state-of-the-art performance in various language tasks.
CodeLlama (Rozière et al., 2023), WizardCoder (Luo et al., 2023), DeepSeek-Coder (AI, 2023): These tokenizers are derived from the tokenizer of Llama 2 (Touvron et al., 2023) which is also trained on the mixed corpus.
SantaCoder (Allal et al., 2023), StarCoder (Li et al., 2023), Replit-code (Repilt, 2023): These tokenizers are specialized for code, having been trained exclusively on programming language datasets, and are thus more adept at handling source code.
CodeGen (Nijkamp et al., 2022), CodeT5 (Wang et al., 2021), CodeT5+ (Wang et al., 2023): Created by Salesforce, these tokenizers are extended based on the vocabulary of GPT2 with additional tokens representing repeating tokens of tabs and white spaces.
Table 2. Percentage of token reduction achieved with SimPy. The “Code” and “Web” in the “Vocab Source” column represent the sources for constructing the tokenizer’s vocabulary: code repositories and internet data, respectively.
{tblr}
cells = c, colsep = 4.5pt, cell11 = r=2, cell12 = r=2, cell13 = r=2, cell14 = c=3, cell25 = c=2, vline1-5,7 = -, hline1,3,17 = -, hline2 = 4-6, Tokenizer & Vocab
Source
Vocab
Size
Tokens
Python SimPy
CodeBert Code 50k 1.33B 0.87B 34.7$\downarrow$
GPT2 Web 50k 1.33B 0.87B 34.7$\downarrow$
CodeLlama Web 32k 0.97B 0.84B 13.5$\downarrow$
WizardCoder Web 32k 0.97B 0.84B 13.5$\downarrow$
DeepSeek-Coder Web 32k 0.97B 0.84B 12.9$\downarrow$
CodeGen Web 51k 0.93B 0.82B 12.6$\downarrow$
CodeT5+ Web 51k 0.93B 0.82B 12.6$\downarrow$
Codex Web 51k 0.93B 0.82B 12.6$\downarrow$
CodeT5 Code 32k 0.91B 0.78B 13.8$\downarrow$
StarCoder Code 49k 0.83B 0.76B 8.6$\downarrow$
SantaCoder Code 49k 0.83B 0.76B 8.8$\downarrow$
Replit-code Code 33k 0.82B 0.75B 8.6$\downarrow$
GPT-3.5 Web 100k 0.71B 0.63B 10.4$\downarrow$
GPT-4 Web 100k 0.71B 0.63B 10.4$\downarrow$
3.2.2. Results
To answer RQ1, we conducted an evaluation involving the representation of code files from our Python dataset in both its original grammar and in SimPy, followed by the tokenization using the same tokenizer for each representation. We created the SimPy dataset by converting the Python dataset with our converter. In tokenizing the SimPy code, we modify the tokenizers to include tokens of SimPy in their vocabularies. In total, 14 tokenizers from popular LLMs are evaluated in our experiments, where each tokenizer’s vocabulary source and size are also documented to offer a comprehensive view of SimPy’s performance across different models. By examining the variation in token numbers, we evaluated SimPy’s effectiveness in reducing token size, thus showcasing the potential benefits of AI-oriented syntax.
As revealed in table 2, SimPy can reduce the number of tokens by 8.6% to 34.7%, depending on the tokenizers. The GPT-4 and GPT-3.5 tokenizers, which are already the most efficient in representing Python source code, show a further reduction of 10.4% in token count with SimPy. For tokenizers trained on code corpora, such as Replit-code and StarCoder, SimPy achieved a token reduction ranging from 8.6% to 13.8%. Tokenizers trained on web-based corpora like CodeGen and CodeT5 also exhibited significant reductions, between 12.6% and 13.5%. The most pronounced impact of SimPy is observed with the least efficient tokenizers, CodeBert and GPT-2, where a remarkable 34.7% reduction in token count was achieved. These promising results highlight SimPy’s potential to reduce token count for source code representation. As estimated by OpenAI (Kaplan et al., 2020), the Floating-point operations (FLOPS) required for generating each token during inference can be regarded as being only relevant to the model size when the context size is fixed. Therefore, a reduction in token count can be directly translated to a decrease in FLOPS at a similar level, resulting in faster inference speeds given the fixed computing speed of the device.
Answer to RQ1: AI-oriented grammar, exemplified using SimPy, effectively reduces the number of tokens required for source code representation, with models like GPT-4 benefiting from a 10.4% reduction. Correspondingly, it leads to a speed up and a computing saving during inference at a similar level.
4. Model Training with AI-oriented Grammar (RQ2)
In this section, we aim to answer RQ2: How can AI models understand AI-oriented grammar? We experimentally investigate whether AI models can retain their accuracy when trained with AI-oriented grammar. We describe our training strategies in Section 4.1 and assess their effectiveness on two language models in Section 4.2.
4.1. Training Strategies
Training AI models with AI-oriented grammar is a pivotal step to enable the model to effectively deal with source code in this new format. Despite the efficiency gains demonstrated by SimPy, such training should not compromise the model’s accuracy. To explore the feasibility of such training, we experiment with two different strategies. Next, we introduce the strategies in the experiment, from tokenizer refining to model training.
Tokenizer Refining SimPy introduces 78 new tokens for the tokenizers to recognize. For example, the “def” keyword of the original Python grammar is replaced by a token “¡def_stmt¿”. Given the existing association between the pre-trained model and its tokenizer, completely retraining the tokenizer on SimPy code to optimize token distribution is impractical. Instead, we opt for a more feasible approach: expanding the tokenizer’s vocabulary to include these new tokens. Correspondingly, this modification requires resizing the embedding matrix ([vocab size * embedding size]) and the output layer ([hidden state size * vocab size]) to fit the expended vocab size. This expansion introduces a few new parameters, mainly in the output layer, around 78 * hidden_size parameters. For instance, modifying a CodeGen (Nijkamp et al., 2022) model with a hidden state size of 2048 introduces around 160 thousand new parameters, a negligible increase (less than 0.01%) in the total parameter count. Moreover, the resizing will randomly initialize both the embedding vector for each new token and the weight of the output layer, which will be updated during the model training.
[…]
4.2. Experiments of RQ2
We first present the experimental setup for RQ2, including the models used, evaluation metrics, and implementation details. Then, we report the experimental results and answer the research questions.
4.2.1. Models
We adopt three widely used models in our research community, namely CodeGen-NL, TinyLlama, and Pythia, whose parameter sizes range between 350M and 1.1B. All these models serve as the initial pre-trained model for our experiments. Though these are not the latest state-of-the-art models, they suffice to validate the feasibility of learning AI-oriented grammar like SimPy. We will further discuss the impact of this decision in Section 7.
CodeGen-NL: CodeGen, proposed by Salesfore (Nijkamp et al., 2022), is an open-sourced language model designed for code generation. It undergoes a multi-phase training process on different datasets, where the model is first trained with natural language datasets and then code datasets. Our experiments utilize its natural language version (CodeGen-350M-nl), produced after the initial phase of its training process, as the foundation model to conduct our experiments.
TinyLlama: TinyLlama (Zhang et al., 2024) is a compact 1.1B language model pre-trained on around 3 trillion tokens, building on the architecture and tokenizer of Llama 2 (Touvron et al., 2023). It shows competitive performance compared to existing open-source language models of similar sizes.
Pythia: Pythia (Biderman et al., 2023) is a suite of LLMs ranging in size from 70M to 12B parameters. It is intended to be used as the baseline for research studies and thus is designed close to currently accepted common practices. Considering the capacity of our computing resources, we use its 1B version.
4.2.2. Evaluation Metrics
We evaluate the model’s performance on the code generation task with the Pass@$k$ metric on HumanEval. To compute Pass@$k$, $k$ code samples are generated for each problem, and a problem is considered solved if any of the $k$ samples pass the unit tests. We report the fraction of problems being successfully solved. The HumanEval dataset, developed by OpenAI (Chen et al., 2021), comprises 164 programming problems, each with a function signature, a docstring, and multiple test cases. Given the function signature and docstring, the model is required to generate the code, which is then tested by executing the test cases. Notably, the function signatures are written using Python’s original grammar. When evaluating the model adapted to SimPy, we convert the function signature into SimPy using the code converter. Similarly, the model-generated SimPy code is subsequently converted into Python to run test cases since the existing testing framework is implemented for Python source code.
4.2.3. Implementation Details
In our experiments, we use the Huggingface Transformers library (Wolf et al., 2020) with Pytorch to implement the models. The experiments of CodeGen-NL are performed on a machine with 48 vCPUs, 512GB RAM, and four RTX A5000 GPUs (24GB RAM), while the other two models are trained on a machine with 28 vCPUs, 200GB RAM, and two RTX A6000 GPUs (48GB RAM). The hyper-parameters of the training are set referring to CodeGen’s hyper-parameters: 8 batch size, 1.8e-4 learning rate, 0.1 weight decay, and 512 context length. During the inference for evaluation, we set the temperature to 0.2 and the top-p to 0.95.
4.2.4. Results
[…]
{tblr}
cells = c, cell21 = r=6, cell81 = r=6, cell141 = r=6, vlines, hline1-2,8,14,20 = -, hline3-4,9-10,15-16 = 2-4, Model & Training Strategy Pass@1 Pass@10
CodeGen-NL  Python 4.51% 7.32%
 100% SimPy 2.93% 5.49%
 Python $\rightarrow$ 10% SimPy 3.11% 3.66%
 Python $\rightarrow$ 20% SimPy 3.66% 4.27%
 Python $\rightarrow$ 50% SimPy 3.96% 6.71%
 Python $\rightarrow$ 100% SimPy 4.82% 9.15%
TinyLlama  Python 10.00% 13.41%
 100% SimPy 5.91% 9.76%
 Python $\rightarrow$ 10% SimPy 2.07% 3.66%
 Python $\rightarrow$ 20% SimPy 3.23% 5.49%
 Python $\rightarrow$ 50% SimPy 5.73% 11.59%
 Python $\rightarrow$ 100% SimPy 10.12% 14.02%
Pythia  Python 5.79% 9.76%
 100% SimPy 7.01% 9.15%
 Python $\rightarrow$ 10% SimPy 1.89% 2.44%
 Python $\rightarrow$ 20% SimPy 3.11% 4.27%
 Python $\rightarrow$ 50% SimPy 4.21% 7.32%
 Python $\rightarrow$ 100% SimPy 5.67% 10.00%
[…]
We report the results in Table 3. Notably, the models trained with SimPy lag behind the Python baseline in terms of accuracy. For example, the Pass@1 and Pass@10 of CodeGen (SimPy) are respectively 2.93% and 5.49%, lower than the ones of CodeGen (Python), which are 4.51% and 7.32%. This could be attributed to SimPy’s limited expressiveness, constraining the models from leveraging knowledge acquired from natural language datasets during pre-training. Consequently, direct training with AI-oriented grammar appears to be an impractical approach.
[…]
Answer to RQ2: AI models, when initially trained with the original grammar and then the AI-oriented grammar, can successfully learn the AI-oriented grammar, retaining or even enhancing their accuracy. For instance, the CodeGen model, originally trained with Python and achieving a 7.32% Pass@10, improved to a 9.15% Pass@10 after the additional training with SimPy.
5. Usage Scenario of AI-oriented Grammar (RQ3)
In this section, we address RQ3: How can AI-oriented grammar support real-world scenarios? We first demonstrate the basic application scenario of AI-oriented grammar, and subsequently, we introduce a novel inference framework designed to broaden the applicability of AI-oriented grammar, followed by an evaluation of the framework’s additional latency.
Figure 4. LEFT: the workflow of the basic usage scenarios of AI-oriented grammar. RIGHT: the workflow of the extended usage scenarios of AI-oriented grammar under DualCode, where the code executor of the AI system in the figure is not necessary.
5.1. Basic usage scenario
The source code, when written in AI-oriented grammar, becomes challenging for human interpretation and is therefore not intended for human display. Consequently, the application of AI-oriented grammar is limited to scenarios where human users do not have access to the generated code. A typical scenario is the AI agents, such as AutoGPT (Significant Gravitas, [n. d.]) and LangChain (Chase, 2022), for regular users rather than developers. For instance, an AI agent tasked with data collection from a website would generate the required crawler script, execute it to gather data, and present the outcomes to the user. End users generally care more about the results than understanding the underlying script since they lack programming knowledge. Therefore, even without additional enhancement, models trained with AI-oriented grammar can be effectively utilized in real-world scenarios. We demonstrate this scenario on the left of Figure 4. In this scenario, an AI-oriented code generated by the model can be executed in two ways: 1) being translated into human-centric code and then executed by its executor; 2) directly being executed by a specific executor for the AI-oriented grammar. Notably, implementing an executor specifically for AI-oriented grammar demands only lightweight engineering efforts as the AI-oriented grammar and its original grammar differ only at the syntax level. Thus, the second method offers a more efficient solution.
5.2. Extended usage scenario
Despite the effectiveness of AI-oriented grammar in certain contexts, many code generation scenarios still require the involvement of humans, where human-readable code is required. To fill this gap, we propose an inference framework for code generation named DualCode. DualCode enables human users to interact with code in human-centric grammar, while the model still leverages the efficiency of AI-oriented grammar during the inference process. The fundamental concept of DualCode is to convert the code between AI-oriented grammar and the original grammar of the same programming language. To achieve this goal, a rule-based code converter should be employed to convert source code into AI-oriented grammar for model comprehension and, inversely for user readability. Such a converter is feasible since both the AI-oriented grammar and original grammar describe the same AST. The identical AST allows the code written in the two grammars to be equivalently converted into each other based on the grammar rules.
We illustrate the workflow of DualCode on the right of Figure 4. It employs two “gates”: an input converter and an output converter. The input converter translates code written in human-centric grammar into AI-oriented grammar for model processing. Similarly, the output converter reverts AI-generated code into human-readable code for user comprehension. Notably, this environment is only for the code, where other inputs, such as natural language, are unaffected. DualCode is a not complicated framework, enabling the lightweight integration of AI-oriented grammar into existing workflows of AI systems. Though being straightforward, it is proposed and investigated for the first time, bridging the gap between efficient AI-oriented code generation and human readability.
5.3. Experiments of RQ3
Table 4. Comparison of average conversion times between Python and SimPy, and the processing speed of the StarCoder tokenizer, based on Huggingface Tokenizers.
{tblr}
cells = c, colsep = 5pt, cell11 = r=2, cell12 = c=2, cell14 = c=2, vline1-3,5 = 1, vline4,6 = 2, vline1-2,4,6 = 1-7, hline1,3,8 = -, hline2 = 2-5, Token num & Huggingface Converter
Encode Decode To SimPy To Python
[0, 100) 0.2ms 0.1ms 0.2ms 0.2ms
[100, 500) 0.7ms 0.6ms 0.9ms 0.8ms
[500, 2000) 2.4ms 2.2ms 3.4ms 3.1ms
[2000, 5000) 6.7ms 6.4ms 12.2ms 10.8ms
[5000, $\infty$) 23.0ms 23.7ms 75.4ms 57.4ms
Given that the DualCode converter adds extra latency to the inference process, a significant concern arises: excessive latency could render the system impractical for real-world applications. To address the concern, we conduct experiments focusing on the converter’s performance. Specifically, we measure the time taken to convert Python code files into SimPy and then back to Python using the converter. As a reference, we evaluate the processing speed of the StarCoder tokenizer, which is based on the widely acknowledged Huggingface Tokenizers library (Moi and Patry, 2023). For this experiment, we categorized Python code files into five distinct groups, based on their token counts, as follows: [0, 100), [100, 500), [500, 2000), [2000, 5000), and [5000, $\infty$). These token counts are determined using the StarCoder tokenizer (Li et al., 2023) on the Python code. We calculate the average processing time for each group, both for the converter and the StarCoder tokenizer.
The findings, presented in Table 4, indicate that the converter’s speed is comparable to that of Huggingface Tokenizers. For code files with fewer than 100 tokens, the converter’s processing time for each conversion is a mere 0.2 ms, only 0.1 ms slower than the Huggingface Tokenizers. For files containing 100 to 500 tokens, the conversion is completed within 1.0 ms. This is not a significant concern, given that over 95% of the dataset’s code files (sourced from real-world repositories) are within the 5000-token range. Therefore, we deduce that the latency induced by the converter is acceptably minimal in most practical scenarios.
Answer to RQ3: Beyond the basic scenarios where human interaction is not required, the application of AI-oriented grammar can be substantially extended by incorporating the DualCode framework. DualCode enables humans to continue using human-centric grammar while AI models leverage the efficiency of AI-oriented grammar. Notably, it imposes negligible latency (under 1 ms for code up to 500 tokens).
6. Related Work
Program Simplification Program simplification has emerged as a valuable approach to enhance the efficiency of code models (Huang et al., 2023; Rabin et al., 2022, 2021; Shi et al., 2023; Bui et al., 2019; Yang et al., 2024). This approach typically involves the elimination of less critical code tokens to streamline model processing. For example, DietCode (Zhang et al., 2022) removes the code tokens that receive the fewest attention weights by CodeBert. Sivand (Rabin et al., 2022) and P2IM (Zheng et al., 2020) simplify the input code according to the outputs of a supplementary model. While these methods considerably boost efficiency, they unavoidably compromise accuracy due to the removal of certain code elements. In contrast, models with AI-oriented grammar, though perhaps less efficient, are able to preserve or even improve accuracy. Most importantly, existing simplification techniques are irreversible, limiting their application to code understanding tasks like summarization and retrieval, rather than code generation. Conversely, code in AI-oriented grammar can be effortlessly reverted to its original form, thus suitable for various code-related tasks.
Tokenization of Source Code Modern LLMs usually preprocess textual datasets using an open-vocabulary tokenization method, Byte-Pair Encoding (BPE) (Sennrich et al., 2015). BPE tokenizes text into subwords based on their frequency in the text corpus, offering a balance between the granularity of tokens and vocabulary breadth. Karampatsis et al. (Karampatsis et al., 2020) first identify the effectiveness of BPE on source code. CodeT5 reveals that BPE trained on source code corpus can reduce over 30% of tokens for code generation, compared with the one trained on natural language corpus. Subsequently, all major LLMs for code generation, such as CodeBERT (Feng et al., 2020), CodeT5 (Wang et al., 2021), SantaCoder (Allal et al., 2023), StarCoder (Li et al., 2023) and CodeLlama (Rozière et al., 2023), adopt BPE as the tokenization method. Further enhancements to BPE for source code have been proposed. For example, Chirkova (Chirkova and Troshin, 2023) suggests that clustering punctuation characters into single tokens can reduce average token length by 17% without impacting model performance. Notably, even though the tokenizers are optimized for source code, they still need to deal with the unnecessary tokens introduced by the human-centric grammar. AI-oriented grammar optimizes the representation of source code in a more fundamental way, which is orthogonal to these existing tokenization methods.
7. Threats to Validity
Constrained Model Selection Our experimental scope in RQ2 is restricted by our computational resources, limiting our evaluation to models with around 1B parameters, e.g., Pythia and TinyLlama. These models are relatively modest in scale. However, while the model size is expanding, the fundamental issue of computation waste caused by human-centric code grammar remains unaddressed. Therefore, the insights derived from our experiments with smaller models are still highly relevant for understanding inefficiency issues in larger models.
Limited Programming Language Our research primarily investigates the implementation of AI-oriented grammar in Python, a language widely utilized by existing LLMs for programming tasks. This initial exploration has shown that AI-oriented grammar effectively reduces computational costs during inference. However, the conclusions drawn from Python may not generalize to other programming languages. We thus leave the exploration of its implementation in other languages as future work.
Inefficient Implementation We implement a proof-of-concept converter to convert the code between SimPy and Python. While this converter provides seamless translation, its efficiency is not optimized. For instance, it is developed in Python, which is less efficient compared to languages like C++. This aspect could potentially result in an underestimation of the converter’s performance in our experimental evaluations.
8. Discussion
Future directions of AI-oriented grammar Our study concentrates on the inefficiencies arising from the mismatch between AI coders and human-centric code grammars. This initial investigation into AI-oriented grammar highlights its feasibility and potential benefits. As an emerging field, AI-oriented grammar presents numerous unexplored questions. For example, an interesting finding from our experiments is that models trained with AI-oriented grammar can even improve the model’s accuracy in code generation tasks. This emphasizes the critical role of grammar as a foundational element for LLMs in grasping code semantics. Designing grammars that are inherently more comprehensible to AI models could significantly enhance their performance. Our current research provides a preliminary insight into this aspect, opening doors for in-depth future studies. Additionally, the process of simplifying grammar, as exemplified by our manual creation of SimPy, raises the question of whether an automated approach could create optimal grammar rules for AI models. A potential solution for simplifying the grammar could be iteratively searching for grammar tokens/structures that can be removed with the help of a parser generator. Moreover, saving the training cost for teaching LLMs AI-oriented grammar is also of great practical value, where a more efficient training method for LLMs to learn new programming grammar is urgently needed. We, therefore, call for the software engineering community to engage further with this promising topic, recognizing its potential to revolutionize the field of AI coders.
Limitations in practice Though extending the applicability of AI-oriented grammar, DualCode relies on a rule-based converter. The converter, we implemented for SimPy, is AST-based, which implicitly requires the input and output code of the models under the DualCode framework to satisfy the grammar correctness. For the output, grammar correctness is a fundamental expectation for a qualified LLM-based assistant. Thus, this requirement from DualCode is not an additional constraint set to the model but aligns with the goal of a reliable AI service. However, it poses challenges when dealing with user-provided input, which may not always be grammatically correct. It is not a concern to models handling natural-language-to-code tasks. However, the requirement may limit the application of SimPy when some tasks involve partial source code as input, such as LLM-based code completion. Addressing this limitation could involve developing an error-tolerant converter or grammar, which is a crucial direction for future research.
Learning the AI-oriented grammar The learning of AI-oriented grammar could be a tricky task. In our experiments, we demonstrate the effectiveness of fine-tuning AI models with SimPy using the next token prediction task. However, this simple fine-tuning strategy requires a large number of SimPy samples, 100% of the dataset in our experiments. A more efficient adaptation process would significantly enhance the utility of AI-oriented grammar. However, current research on how AI models learn code grammar is still limited. Although studies (Wan et al., 2022; Chen et al., 2022; Ma et al., 2022) have shown that LLMs typically grasp code grammar knowledge in their initial layers, the exact learning mechanism remains unclear. Therefore, a thorough analysis in this area is much needed.
Utility of AI-oriented grammar In this paper, we demonstrate the effectiveness of the sequential training scheme, where the model is initially trained with the original grammar and then the AI-oriented grammar. It achieves an equivalent, or even improved, performance compared to the model trained merely with the original grammar. Such a training method incurs an increase in the cost of the model training. For example, training CodeGen on the original Python dataset costs 183,628 training steps, and 100,288 additional steps are taken during the further finetuning on the 100% SimPy dataset. Nevertheless, mastering AI-oriented grammar still reduces energy consumption in the long run. Training is performed only once or occasionally, while inference tasks can be continuous and massive after the system is deployed. The post-deployment operational cost is a primary component of the overall cost, sometimes reaching 90% of total expenses (Desislavov et al., 2023). Consequently, despite the additional costs incurred during training, implementing AI-oriented grammar remains highly beneficial from a practical standpoint.
9. Conclusion and Future work
In this paper, we, for the first time, propose the concept of AI-oriented grammar to address the inefficiency of AI coders in processing the code written in human-centric grammar. Through an empirical study guided by three research questions, we successfully demonstrate the feasibility and potential of this novel concept. During our research, we have developed the first-ever AI-oriented grammar for Python. Additionally, we introduced an inference framework designed to empower models to effectively process both AI-oriented and human-centric grammars within a programming language. In the future, our focus will shift towards investigating how AI-oriented grammar can further refine the accuracy of AI models.